# Mixing Time Bounds for the Gibbs Sampler under Isoperimetry

### v1

@@ Acknowledgments
Alexander Goyal is fully funded by the Roth scholarship. George Deligiannidis is supported by the Engineering and Physical Sciences Research Council [grant number EP/Y018273/1].
@@ Appendix A Appendix
Define the function $N_{q}(x)=\left\lvert x\right\rvert^{q}\cdot\log(1+\left\lvert x\right\rvert^{q})$ for all $x\in\mathbb{R}^{d}$. Given a function $f:\mathbb{R}^{d}\to\mathbb{R}$, we denote its Orlicz norm associated to $N_{q}$ by
@@
Let $\left\|f\right\|_{q}$ also denote the $L^{q}_{\pi}$-norm of $f$, as defined in Subsection 4.1. We now introduce two functional inequalities in the Orlicz space. Let $\mathtt{c}_{q}>0$ and $\tilde{\mathtt{c}}_{q}>0$ denote the largest constants such that, for all locally Lipschitz functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
$\mathtt{c}_{q}\left\|f-\mathbb{E}[f]\right\|_{N_{q}}^{q}$ $\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right],$ (17)
$\tilde{\mathtt{c}}_{q}\left\|f-m[f]\right\|_{N_{q}}^{q}$ $\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right],$ (18)
where $m[f]$ denotes a median of $f$. Bobkov and Zegarlinski (2005, Proposition 3.1) showed that $\mathrm{LSI}_{q}$ is equivalent to the functional inequality (17), with constants satisfying $\mathtt{c}_{q}/16\leq\mathtt{s}_{q}\leq 7\mathtt{c}_{q}$. Moreover, by Milman (2009a, Lemma 2.1), we have $\tilde{\mathtt{c}}_{q}/2\leq\mathtt{c}_{q}\leq 3\tilde{\mathtt{c}}_{q}$. The functional inequality (18) is therefore equivalent to $\mathrm{LSI}_{q}$, with constants satisfying
$\frac{\tilde{\mathtt{c}}_{q}}{32}\leq\mathtt{s}_{q}\leq 21\tilde{\mathtt{c}}_{q}.$ (19)
@@
${\tilde{\mathtt{c}}_{q}}^{1/q}\left\|g\right\|_{N_{q}}\leq\left\|\left\lvert\nabla g\right\rvert\right\|_{q}.$ (20)
Since the solution to the equation $1/u+1/r=1/q$ is $u=qr/(r-q)$, an application of Hölder’s inequality gives
@@
The last inequality follows from Bobkov and Zegarlinski (2005, Lemma 3.2). Furthermore, for all $\lambda>0$, we have
@@
${\tilde{\mathtt{c}}_{q}}^{1/q}\left\|f\right\|_{N_{r}}^{r/q}\leq\left(\frac{5}{4}\right)^{1/q-1/r}\frac{r}{q}\left\|f\right\|_{N_{r}}^{r/q-1}\cdot\left\|\left\lvert\nabla f\right\rvert\right\|_{r}.$
@@
$\left(\frac{4}{5}\right)^{r/q-1}\left(\frac{q}{r}\right)^{r}{\tilde{\mathtt{c}}_{q}}^{r/q}\left\|f\right\|_{N_{r}}^{r}\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{r}\right].$
@@
$\left(\frac{4}{5}\right)^{r/q-1}\left(\frac{q}{r}\right)^{r}{\tilde{\mathtt{c}}_{q}}^{r/q}\leq\tilde{\mathtt{c}}_{r}.$
@@
Let $S$ $\subseteq\mathbb{R}^{d}$ with $\pi(S)\leq 1/2$. Note that in this case, $\pi(S^{c})\geq\pi(S)$. Define the sets
$S_{1}$ =$=\left\{x\in S:P(x,S^{c})\leq\frac{\varepsilon}{2}\right\},$
$S_{2}$ =$=\left\{x\in S^{c}:P(x,S)\leq\frac{\varepsilon}{2}\right\},$
@@
By the contrapositive of the $(\delta,\varepsilon)$-close coupling condition, this implies $\left\lvert x-y\right\rvert>\delta$, and consequently, $\mathcal{D}(S_{1},S_{2})\geq\delta$. This leads to
@@
The derivative of $\rho$ is given by $\rho^{\prime}(x)=\log(1/x)-1$, which is positive for $x<e^{-1}$ and negative for $x\geq e^{-1}$. This establishes the first part of the lemma. Furthermore, for $e^{-1}\leq x\leq 1$, observe that $\rho^{\prime}(x)\in[-1,0]$, implying that $\rho$ is Lipschitz continuous on this interval with constant 1.
To prove the second part, we consider the following cases. First, if $x+h<e^{-1}$, then both $x$ and $x+h$ lie in the region where $\rho$ is increasing. Therefore, $\rho(x+h)\geq\rho(x)\geq\rho(x)-h$, as required. Next, consider the case where $x+h\geq e^{-1}$. If $x\geq e^{-1}$, since $\rho$ is decreasing and Lipschitz continuous on this region, we have as desired
$\rho(x)-\rho(x+h)=\left\lvert\rho(x)-\rho(x+h)\right\rvert\leq h.$
If $x<e^{-1}$, we split the integral representation of $\rho(x+h)-\rho(x)$ as follows:
$\rho(x+h)$ $=\rho(x)+\int_{x}^{e^{-1}}\rho^{\prime}(s)\mathrm{d}s+\int_{e^{-1}}^{x+h}\rho^{\prime}(s)\mathrm{d}s$
$\geq\rho(x)+\int_{e^{-1}}^{x+h}\rho^{\prime}(s)\mathrm{d}s$
$\geq\rho(x)-\left(x+h-e^{-1}\right)$
$\geq\rho(x)-h.$
This establishes the second part of the lemma. ∎
@@
Let $T$ be the random variable representing the number of i.i.d. draws from the distribution $\mathrm{Unif}[d]$ required to observe all $d$ unique values. Following Mitzenmacher and Upfal (2017, Subsection 2.4.1), we have
$\mathbb{E}[T]=d\sum_{j=1}^{d}\frac{1}{j}.$
Approximating the above sum with the logarithm, we have
$d\cdot\log d\leq\mathbb{E}[T]\leq d\cdot\log d+d.$
An application of Markov’s inequality finally yields
$\mathbb{P}(\Xi=[d])$ $=\mathbb{P}(T\leq N)$
$\geq\mathbb{P}(T\leq 4d\cdot\log d)$
$\geq\mathbb{P}(T\leq 2(d\cdot\log d+d))$
$=1-\mathbb{P}(T>2(d\cdot\log d+d))$
$\geq 1-\frac{\mathbb{E}[T]}{2(d\cdot\log d+d)}$
$\geq\frac{1}{2}.$
@@
Consider a partition of the space such that $\pi=\pi\left(\theta,s\right)$ with $\theta\in\mathbb{R}^{d_{1}}$, $s\in\mathbb{R}^{d_{2}}$, where $d_{1}+d_{2}=d$. Write $\pi\left(\theta,s\right)\propto e^{-U\left(\theta,s\right)}$. The marginal distribution of $\pi$ with respect to $s$ can be expressed as $\pi\left(s\right)\propto e^{-V\left(s\right)}$, where
@@
Additionally, the conditional distribution of $\pi$ is
$\pi(\mathrm{d}\theta\mid s)=\frac{e^{-U(\theta,s)}\mathrm{d}\theta}{\int e^{-U(\eta,s)}\mathrm{d}\eta}.$
If $U$ is Lipschitz continuous, then the gradient $\nabla U$ exists almost everywhere and satisfies $\left\lvert U(\theta,s)\right\rvert\leq L$. The gradient of $V$ can be written as
@@
where $\nabla_{s}U(\theta,s)$ represents the gradient of $U(\theta,s)$ with respect to the coordinates of $s$. Therefore,
$\left\lvert\nabla V(s)\right\rvert=\int\left\lvert\nabla_{s}U(\theta,s)\right\rvert\pi(\mathrm{d}\theta\mid s)\leq L$
almost everywhere, which shows that $V$ is also $L$-Lipschitz continuous.
Now suppose that $U$ is $L$-smooth, meaning that the Hessian $\nabla^{2}U$ exists almost everywhere and satisfies $\nabla^{2}U(\theta,s)\preceq LI_{d}$, where $I_{d}$ denotes the $d$-dimensional identity matrix. Writing $w^{2}=ww^{T}$ for a vector $w\in\mathbb{R}^{d_{2}}$, the Hessian of $V$ is
$\nabla^{2}V(s)=\int\nabla^{2}U(\theta,s)\pi(\mathrm{d}\theta\mid s)+\left(\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\right)^{2}-\int\left(\nabla_{s}U(\theta,s)\right)^{2}\pi(\mathrm{d}\theta\mid s).$
By the Cauchy-Schwarz inequality,
$\left(\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\right)^{2}-\int\left(\nabla_{s}U(\theta,s)\right)^{2}\pi(\mathrm{d}\theta\mid s)\preceq 0.$
Thus,
$\nabla^{2}V(s)\preceq\int\nabla^{2}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\preceq LI_{d_{2}},$
implying that $V$ is also $L$-smooth. ∎

### v2

@@ Acknowledgments
AG was fully funded by the Roth Scholarship provided by the Department of Mathematics, Imperial College London. GD was supported during a large part of this work by the Engineering and Physical Sciences Research Council [grant number EP/Y018273/1].
@@ Appendix A Appendix
Define the function $N_{q}(x)=\left\lvert x\right\rvert^{q}\cdot\log(1+\left\lvert x\right\rvert^{q})$ for all $x\in\mathbb{R}^{d}$. Given a function $f:\mathbb{R}^{d}\to\mathbb{R}$, denote its Orlicz norm associated to $N_{q}$ by
@@
Let $\left\|f\right\|_{q}$ also denote the $L^{q}_{\pi}$-norm of $f$, as defined in Subsection 4.1. We now introduce two functional inequalities in the Orlicz space. Let $\mathtt{k}_{q}>0$ and $\tilde{\mathtt{k}}_{q}>0$ denote the largest constants such that, for all locally Lipschitz functions $f:\mathbb{R}^{d}\to\mathbb{R}$,
$\mathtt{k}_{q}\left\|f-\mathbb{E}[f]\right\|_{N_{q}}^{q}$ $\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right],$ (21)
$\tilde{\mathtt{k}}_{q}\left\|f-m[f]\right\|_{N_{q}}^{q}$ $\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{q}\right],$ (22)
where $m[f]$ denotes a median of $f$. Bobkov and Zegarlinski (2005, Proposition 3.1) showed that $\mathrm{LSI}_{q}$ is equivalent to the functional inequality (21), with constants satisfying $\mathtt{k}_{q}/16\leq\mathtt{s}_{q}\leq 7\mathtt{k}_{q}$. Moreover, by Milman (2009a, Lemma 2.1), $\tilde{\mathtt{k}}_{q}/2\leq\mathtt{k}_{q}\leq 3\tilde{\mathtt{k}}_{q}$. The functional inequality (22) is therefore equivalent to $\mathrm{LSI}_{q}$, with constants satisfying
$\frac{\tilde{\mathtt{k}}_{q}}{32}\leq\mathtt{s}_{q}\leq 21\tilde{\mathtt{k}}_{q}.$ (23)
@@
${\tilde{\mathtt{k}}_{q}}^{1/q}\left\|g\right\|_{N_{q}}\leq\left\|\left\lvert\nabla g\right\rvert\right\|_{q}.$ (24)
Since the solution to the equation $1/u+1/r=1/q$ is u=qr/(rq)u=qr/(r-q), an application of Hölder’s inequality gives
@@
The last inequality follows from Bobkov and Zegarlinski (2005, Lemma 3.2). Furthermore, for all $\lambda>0$,
@@
${\tilde{\mathtt{k}}_{q}}^{1/q}\left\|f\right\|_{N_{r}}^{r/q}\leq\left(\frac{5}{4}\right)^{1/q-1/r}\frac{r}{q}\left\|f\right\|_{N_{r}}^{r/q-1}\cdot\left\|\left\lvert\nabla f\right\rvert\right\|_{r}.$
@@
$\left(\frac{4}{5}\right)^{r/q-1}\left(\frac{q}{r}\right)^{r}{\tilde{\mathtt{k}}_{q}}^{r/q}\left\|f\right\|_{N_{r}}^{r}\leq\mathbb{E}\left[\left\lvert\nabla f\right\rvert^{r}\right].$
@@
$\left(\frac{4}{5}\right)^{r/q-1}\left(\frac{q}{r}\right)^{r}{\tilde{\mathtt{k}}_{q}}^{r/q}\leq\tilde{\mathtt{k}}_{r}.$
@@
Let $S$ $\subseteq\mathbb{R}^{d}$ with $0<\pi(S)\leq 1/2$. Note that in this case, $\pi(S^{c})\geq\pi(S)$. Define the sets
$S_{1}$ $=\left\{x\in S:P(x,S^{c})\leq\frac{\varepsilon}{2}\right\},$
$S_{2}$ $=\left\{x\in S^{c}:P(x,S)\leq\frac{\varepsilon}{2}\right\},$
@@
By the contrapositive of the $(\delta,\varepsilon)$-close coupling condition, this implies that $\left\lvert x-y\right\rvert>\delta$, and consequently, $\mathcal{D}(S_{1},S_{2})\geq\delta$. This leads to
@@
Since $\rho(0)=0$ and $\log(1/x)\geq 0$ for $x\in(0,1]$, the function $\rho$ is positive. The first and second derivatives of $\rho$ are given by
$\rho^{\prime}(x)=\log\left(\frac{1}{x}\right)-1,\qquad\rho^{\prime\prime}(x)=-\frac{1}{x}.$
Since $\rho^{\prime\prime}$ is negative on its domain, the function $\rho$ is concave. Moreover, $\rho^{\prime}(x)=0$ at $x=e^{-1}$, and hence $\rho$ attains its maximum at $x=e^{-1}$. This proves the first part of the lemma. For the second part, by concavity of $\rho$, a Taylor expansion at $x+h$ yields
$\rho(x)\leq\rho(x+h)-h\rho^{\prime}(x+h).$
Since $\rho^{\prime}(x+h)\geq-1$, it follows that $\rho(x)\leq\rho(x+h)+h$, and therefore
$\rho(x)-h\leq\rho(x+h).$
@@
For any $k\in[d]$, the probability that $k$ is not drawn is
$\mathbb{P}(k\notin\Xi)=\left(1-\frac{1}{d}\right)^{N}\leq e^{-N/d}.$
By a union bound,
$\mathbb{P}(\Xi=[d])$ $=1-\mathbb{P}(\Xi\neq[d])$
$=1-\mathbb{P}\left(\bigcup_{k=1}^{d}\{k\notin\Xi\}\right)$
$\geq 1-\sum_{k=1}^{d}\mathbb{P}(k\notin\Xi)$
$\geq 1-de^{-N/d}.$
In particular, if $N=\lceil 2d\cdot\log d\rceil$,
$\mathbb{P}(\Xi=[d])\geq 1-de^{-2\cdot\log d}=1-\frac{1}{d}.$
Since we assume $d\geq 2$, it follows that $\mathbb{P}(\Xi=[d])\geq 1/2$. ∎
@@
Consider a partition of space such that $\pi=\pi(\theta,s)$ with $\theta\in\mathbb{R}^{d_{1}}$, $s\in\mathbb{R}^{d_{2}}$, where $d_{1}+d_{2}=d$. Write $\pi(\theta,s)\propto\exp(-U(\theta,s))$. If $U$ is $L$-Lipschitz continuous, then the gradient $\nabla U$ exists almost everywhere and satisfies $\left\lvert\nabla U(\theta,s)\right\rvert\leq L$. The marginal distribution of $s$ can be expressed as $\pi(s)\propto\exp(-V(s))$, where
@@
Additionally, the conditional distribution given $s$ is
$\pi(\mathrm{d}\theta\mid s)=\frac{e^{-U(\theta,s)}\mathrm{d}\theta}{\int e^{-U(\psi,s)}\mathrm{d}\psi}.$
In particular, the gradient of $V$ can be written as
@@
where $\nabla_{s}U(\theta,s)$ represents the gradient of $U(\theta,s)$ with respect to the coordinates of $s$. Consequently,
$\left\lvert\nabla V(s)\right\rvert\leq\int\left\lvert\nabla_{s}U(\theta,s)\right\rvert\pi(\mathrm{d}\theta\mid s)\leq L,$
which shows that $V$ is also $L$-Lipschitz continuous.
Suppose now that $\theta\in\mathbb{R}$ and $s\in\mathbb{R}^{d-1}$. If $U$ is $L$-smooth, then the Hessian matrix $\nabla^{2}U$ exists almost everywhere and its spectral norm satisfies
$\left\|\nabla^{2}U(\theta,s)\right\|_{\mathrm{spec}}=\sup\left\{\frac{x^{T}\nabla^{2}U(\theta,s)y}{\left\lvert x\right\rvert\left\lvert y\right\rvert}:x,y\in\mathbb{R}^{d},\,x\neq 0,\,y\neq 0\right\}\leq L.$ (25)
Let $z\in\mathbb{R}^{d-1}$, $z\neq 0$. Taking $x=(0,z)$ and $y=(1,0)$ in (25) gives
$\left\lvert\nabla^{2}_{\theta s}U(\theta,s)\right\rvert=\sup\left\{\frac{z^{T}\nabla^{2}_{\theta s}U(\theta,s)}{\left\lvert z\right\rvert}:z\in\mathbb{R}^{d-1},\,z\neq 0\right\}\leq L.$
Moreover, taking $x=y=(0,z)$ and x=y=(0,z)x=-y=(0,z) in (25) yields
$LI_{d-1}\preceq\nabla_{ss}^{2}U(\theta,s)\preceq LI_{d-1}.$
Writing $z^{2}=zz^{T}$ for the rank-one matrix associated with $z\in\mathbb{R}^{d-1}$, the Hessian matrix $\nabla^{2}V$ admits the decomposition
$\nabla^{2}V(s)$ $=\int\nabla^{2}_{ss}U(\theta,s)\pi(\mathrm{d}\theta\mid s)+\left(\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\right)^{2}-\int\left(\nabla_{s}U(\theta,s)\right)^{2}\pi(\mathrm{d}\theta\mid s)$
$=\int\nabla^{2}_{ss}U(\theta,s)\pi(\mathrm{d}\theta\mid s)-\int\left(\nabla_{s}U(\theta,s)-\int\nabla_{s}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\right)^{2}\pi(\mathrm{d}\theta\mid s)$
$=\int\nabla^{2}_{ss}U(\theta,s)\pi(\mathrm{d}\theta\mid s)-\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[\nabla_{s}U(\theta,s)],$ (26)
where $\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[\nabla_{s}U(\theta,s)]$ denotes the covariance matrix of the vector $\nabla_{s}U(\theta,s)$. Since $\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[\nabla_{s}U(\theta,s)]\succeq 0$, it follows from (26) that
$\nabla^{2}V(s)\preceq\int\nabla^{2}_{ss}U(\theta,s)\pi(\mathrm{d}\theta\mid s)\preceq LI_{d-1}.$ (27)
We now derive an upper bound on the covariance matrix $\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[\nabla_{s}U(\theta,s)]$. If each one-dimensional conditional distribution $\pi(\cdot\mid s)$ satisfies $\mathrm{PI}_{2}$ with constant $\mathtt{c}_{2}$, then, for any $z\in\mathbb{R}^{d-1}$,
$z^{T}\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[\nabla_{s}U(\theta,s)]z$ $=\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[z^{T}\nabla_{s}U(\theta,s)]$
$\leq\frac{1}{\mathtt{c}_{2}}\cdot\mathbb{E}_{\theta\sim\pi(\cdot\mid s)}\left[\left(z^{T}\nabla_{\theta s}^{2}U(\theta,s)\right)^{2}\right]$
$\leq\frac{\left\lvert z\right\rvert^{2}}{\mathtt{c}_{2}}\cdot\mathbb{E}_{\theta\sim\pi(\cdot\mid s)}\left[\left\lvert\nabla_{\theta s}^{2}U(\theta,s)\right\rvert^{2}\right]$
$\leq\frac{L^{2}\left\lvert z\right\rvert^{2}}{\mathtt{c}_{2}}.$
It follows that
$\mathrm{Var}_{\theta\sim\pi(\cdot\mid s)}[\nabla_{s}U(\theta,s)]\preceq\frac{L^{2}}{\mathtt{c}_{2}}I_{d-1}.$
Substituting this bound into (26), we obtain
$\nabla^{2}V(s)\succeq\int\nabla^{2}_{ss}U(\theta,s)\pi(\mathrm{d}\theta\mid s)-\frac{L^{2}}{\mathtt{c}_{2}}I_{d-1}\succeq-\left(L+\frac{L^{2}}{\mathtt{c}_{2}}\right)I_{d-1}.$ (28)
Combining (27) and (28), we conclude that
$\left(L+\frac{L^{2}}{\mathtt{c}_{2}}\right)I_{d-1}\preceq\nabla^{2}V(s)\preceq LI_{d-1},$
and therefore $V$ is $L^{\prime}$-smooth with
$L^{\prime}=L+\frac{L^{2}}{\mathtt{c}_{2}}.$